# PLAYER*: Enhancing LLM-based Multi-Agent Communication and Interaction in Murder Mystery Games

### v2

@@ 4.3 Ablation Studies
We compared agents’ behavior across different numbers of rounds and with varying numbers of questions agents can ask per round, as shown in Figure 7. Performance peaks around round 3, after which it shows variability, with some rounds experiencing slight declines or plateauing in scores despite more rounds or questions. This indicates that after a rapid initial learning or adaptation phase, where agents effectively use additional questions to enhance their understanding and strategies, the value of information gained from conversations tends to converge. These results also provide empirical support for the assumption made in our methodology that the more inquiries we pose to an agent, the expected reward associated with questioning the same agent decreases. For the main results we reported, we use the original setting for number of rounds in MMG, which is 3, and based on the outcomes of the ablation studies, we chose the most effective number of questions to ask each round, which is $1$ question.
@@ 5 Related Works
Multi-agent reinforcement learning marking significant progress in complex games (Lanctot et al., 2017; Perolat et al., 2022; Bakhtin et al., 2023). However, these methods often require extensive time and computational resources and lack linguistic communication capabilities. With the emergence of LLMs, there’s a shift of focus towards improving multi-agent language communication, evidenced by advancements in various games and scenarios, such as werewolf (Xu et al., 2023), avalon (Wang et al., 2023; Shi et al., 2024), interactive narrative (Zhao et al., 2024b), MMGs (Wu et al., 2024), and survival games (Toy et al., 2024). Exemplified by AlphaGo (Silver et al., 2017), self-play learning frameworks (Fu et al., 2023; Chen et al., 2024) are proposed to improve LLMs’ performance. Compared to classic methods (Wang and Shen, 2024), agents based on LLMs are capable of inferencing across a broader range of scenarios (Lin et al., 2023), even with some ability of theory of mind (Zhou et al., 2023) to infer other agent’s mental states. However, they have also been found to inherit biases that limit their inferential abilities (Xie et al., 2023; Chuang et al., 2024). Works have also explored utilising LLMs as the environment (Zhang et al., 2024) or update actions (Zhao et al., 2024a) for agents.
@@
Alignment through human feedback offers more consistent training compared to reinforcement learning (Liang et al., 2024), but obtaining this feedback can be expensive. To address this, approaches like self-instruct (Wang et al., 2022; Liu et al., 2023b), self-reflect (Yao et al., 2023), self-alignment (Sun et al., 2023; Li et al., 2024b), and few-shot planning (Song et al., 2023), have been introduced. These approach was also adapted to search for optimal tools (Du et al., 2024), interact with grounded environments (Ouyang and Li, 2023; Ismail et al., 2024). We were also inspired by stochastic search methods utilised for robots in planning optimal strategies in complex environments (Gammell et al., 2020; Liang et al., 2024), shares many similarities with optimisation tasks for agents (Singh et al., 2023).
@@ 6 Conclusion
PLAYER* addresses limitations of LLM-based agents in complex reasoning using an anytime sampling-based planner with sensors and pruners. Our approach enables efficient, question-driven searching. We propose a quantifiable evaluation method, contribute the WellPlay dataset, and demonstrate PLAYER*’s superiority, advancing effective reasoning agents in complex environments.
Limitations
We identify some of the limitations below:
Limited Annotation Scope
Due to the cost of manual annotation and running experiments, we were unable to annotate all 100 games in the Conan dataset. However, we believe that the current scope of our study is sufficient to demonstrate the efficiency of our methodology. Future research can expand upon our work by annotating additional games and applying our approach to MMGs on a larger scale.
Limited Evaluation of LLMs
Due to the high cost associated with running experiments on LLMs, our experiments were primarily conducted using GPT-3.5. To further support our findings, we also performed smaller-scale experiments on LLaMa, which yielded results consistent with those obtained from GPT-3.5. However, we acknowledge that our inability to run more extensive evaluations across a wider range of popular LLMs, such as GPT-4 or lesser-known models, may have limited our ability to uncover additional insights. The cost of API usage for models like GPT-4 and the substantial running time required for models like LLaMa posed significant constraints on our experimental scope. These limitations highlight the need for future research to explore more cost-effective and efficient methods for evaluating and comparing the performance of various LLMs in the context of complex social deduction games.
@@ A.1 Detailed Rules
In the voting phase, each player needs to cast their vote for who they think is the murderer in each case. If the player with the most votes is the murderer, the civilian players win. Otherwise, the murderer players win.
@@ A.2 Procedure
In this stage, players vote anonymously to determine their suspicious regarding the identity of the murderer. Each player has one vote.
@@ A.3 Example: Solitary Boat Firefly Script
As an illustrative example, we examine the Solitary Boat Firefly murder mystery script, involving six players: [ “Tian Chou", “Zhou Lianyi", “Xi Yan", “Yu Sunian", Yannan", and “Zhou Chitong" ], along with four victims: [ “Zhou Mengdang", “Bao Liu", “Cui Shouheng", and “Wang Xi Rong" ]. We model these characters through a set of agents, denoted as $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$, where $N_{a}=6$ corresponds to the number of players. In parallel with the four victims, we define a set of victims $\mathcal{V}=\{v_{i}\}_{i=1}^{N_{v}}$, where $N_{v}=4$, representing the total number of victims in the scenario.
@@
  • Suspicion State ($s_{1}$): The script features four victims, each harboring suspicions towards the remaining five players. Consequently, $s_{1}$ is represented as a $4\times 5$ matrix, with each row corresponding to a victim and each column reflecting the suspicions they hold against other players. For example, a first row of $[1,0,0,1,0]$ signifies that victim “Cui Shouheng" suspects both “Zhou Lianyi" and “Yannan". So we can get $suspect\_list$ of victim “Cui Shouheng" is [“Zhou Lianyi" ,“Yannan"]
@@
During the game, the dialogue history is recorded in $D$, and as the game progresses, inquiries and responses are conducted in accordance with Algorithm 1, leading to updates in $s_{i}$.
@@ B.1 Scripts
    Background Script. For each character, there is corresponding background script includes all the information from their perspective. For example, for “Sylvia Costa” in the script “Death Wears White”, it is:
@@ D. Father Tom is hate of Drake Li Morette
Table A1: Examples of Each Type of Our Evaluation Questions
@@ 10 Do you have any personal opinions or theories about the case?
Table A2: Predefined questions for Werewolf method.
@@ B.2 Evaluation questions
The evaluation questions are devised based on our annotations and aim to assess the understanding of the intricate relationships and narratives present within the game scripts. These questions are segmented into three categories, with each category of questions and examples shown in Table A1.
@@
Detailed script statistics and evaluation question metrics are presented in Table1.
B.3 Overall Performance Computing
@@
B.4 Annotation Details
B.5 Disclaimers of Risks
Appendix C Implementation
C.1 Implementation Details
RAG (Retrieval-Augmented Generation)
For retrieval enhancement, we utilised the FAISS [^4] https://github.com/facebookresearch/faiss library to build a vector database, creating FAISS indices using the L2 distance. Embeddings were obtained via the Azure API’s text-embedding-ada-002 service. Scripts were stored in segments, with each segment having a maximum length of 50 tokens. For dialog records, a question-and-answer pair was stored as a single segment. During retrieval, the maximum script length and dialog length inserted into the prompt were both set to 4000 tokens. For evaluation, these maximum lengths were increased to 5000 tokens.
Experiment
C.2 Comparison Models
@@
      Selecting questions from a predefined question written by the human specialist, based on the current dialogue and script. The list of expertly devised questions is presented in Table A2.

### v3

@@ 4.7 Ablation Studies
We conducted comprehensive ablation studies to evaluate the effectiveness of the key components of our approach. Removing the Pruner or Sensor modules adversely affected both performance and cost reduction, confirming their utility; results are provided in Table 12. Additionally, we experimented with different sensor configurations and numbers of rounds/questions, results are shown in Figure 7. Our analyses revealed that a carefully chosen subset of three sensors strikes a balance between performance gains and computational costs and that around three rounds of interaction with one question per round is near-optimal. These experimental setups, alongside full performance comparisons and in-depth evaluations, are presented in the Appendix D.3.
@@ 5 Related Works
Multi-agent reinforcement learning marking significant progress in complex games (Lanctot et al., 2017; Perolat et al., 2022; Bakhtin et al., 2023). However, these methods often require extensive computational resources and lack linguistic communication capabilities. With the emergence of LLMs, there’s a shift of focus towards improving multi-agent language communication, evidenced by advancements in various games and scenarios, such as werewolf (Xu et al., 2023a), avalon (Wang et al., 2023; Shi et al., 2024), interactive narrative (Zhao et al., 2024b), MMGs (Wu et al., 2024), and survival games (Toy et al., 2024). Exemplified by AlphaGo (Silver et al., 2017), self-play learning frameworks (Fu et al., 2023; Chen et al., 2024) are proposed to improve LLMs’ performance. Compared to classic methods (Wang and Shen, 2024), LLMs based methods can perform inference across a wider variety of scenarios (Lin et al., 2023), even with some ability of theory of mind (Zhou et al., 2023) to infer other agent’s mental states. However, the inherit biases that can potentially limit their inferential abilities (Xie et al., 2023; Chuang et al., 2024; Wang et al., 2024). Works have also explored utilising LLMs as the environment (Zhang et al., 2024) or update actions (Zhao et al., 2024a) for agents.
@@
Alignment through human feedback offers more consistent training compared to reinforcement learning (Liang et al., 2024), but obtaining this feedback can be expensive. Therefore, approaches like self-instruct (Wang et al., 2022; Liu et al., 2023), self-reflect (Yao et al., 2023), self-alignment (Sun et al., 2023; Li et al., 2024b), and few-shot planning (Song et al., 2023), have been introduced. These approaches was also adapted to search for optimal tools (Du et al., 2024), and interact with grounded environments (Ouyang and Li, 2023; Ismail et al., 2024). We were also inspired by stochastic search methods for robots in planning optimal strategies in complex environments (Gammell et al., 2020; Liang et al., 2024), shares many similarities with optimisation tasks for agents (Singh et al., 2023).
@@ 6 Conclusion
This study introduces Questum, a framework that tackles the challenges of LLM-driven agents in MMGs through sensor-based modeling, information gain-driven questioning, and suspect pruning. With the WellPlay dataset enabling robust evaluation, Questum  demonstrates superior reasoning, efficiency, and human-like interaction quality. These advancements pave the way for more effective and engaging AI systems in dynamic multi-agent scenarios.
Limitation
Annotation Scope
Murder Mystery Games are built on intricately written human-authored stories, featuring dense plots, complex character relationships, and unexpected twists to enhance gameplay. These characteristics make annotation particularly challenging, explaining the scarcity of labeled datasets for MMGs. Existing resources include only 56 inferential questions across 4 selected MMGs.
While this dataset is still insufficient for training due to the task’s inherent complexity—and the impracticality of creating a large dataset for such scenarios—we believe it adequately demonstrates the effectiveness of our methodology. Future research could explore synthetic data generation with enhanced quality control to further expand the dataset.
Generalisation Across Domains
While the proposed framework demonstrates strong performance for MMGs, its generalisability to other multi-agent or language-driven tasks has not been fully explored. Applying this approach to domains with different interaction dynamics or less structured narratives may require substantial modifications.
Future research could extend the framework to other multi-agent scenarios, such as collaborative storytelling, negotiation games, or real-world applications like customer service. Domain adaptation techniques could help refine the framework for broader use cases.
Resource Dependence
The framework depends on powerful LLMs for sensor-based reasoning and question generation, which can be resource-intensive. Scaling this approach to larger or real-time systems may require additional computational and memory resources, limiting accessibility for smaller-scale applications.
Efforts to optimise computational efficiency, such as caching mechanisms, or lightweight models for real-time applications, could make the framework more accessible and scalable.
@@ A.1 Detailed Rules
In the voting phase, each player needs to cast their vote for who they think is the murderer in each case. If the player with over 50% of the votes is the murderer, the civilian players win. Otherwise, the murderer players win.
@@ A.2 Procedure
In this stage, players vote anonymously to determine their suspicions regarding the identity of the murderer. Each player has one vote.
@@ A.3 Example: Solitary Boat Firefly Script
As an illustrative example, we examine the Solitary Boat Firefly murder mystery script, involving six players: [ “Tian Chou", “Zhou Lianyi", “Xi Yan", “Yu Sunian", Yannan", and “Zhou Chitong" ], along with four victims: [ “Zhou Mengdang", “Bao Liu", “Cui Shouheng", and “Wang Xi Rong" ]. We model these characters through a set of agents, denoted as $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$, where $N_{a}=6$ corresponds to the number of players. In parallel with the four victims, we define a set of victims $\mathcal{V}=\{v_{k}\}_{k=1}^{N_{v}}$, where $N_{v}=4$, representing the total number of victims in the scenario.
@@
  • Suspicion State ($s_{i}jk$): The script features four victims, each harbouring suspicions towards the remaining five players. Consequently, $s_{i}$ is represented as a $4\times 5$ matrix, with each row corresponding to a victim and each column reflecting the suspicions they hold against other players. For example, a first row of $[1,0,0,1,0]$ signifies that victim “Cui Shouheng" suspects both “Zhou Lianyi" and “Yannan". So we can get $suspect\_list$ of victim “Cui Shouheng" is [“Zhou Lianyi" ,“Yannan"]
@@ B.1 Scripts
    Background Script. For each character, there is a corresponding background script includes all the information from their perspective. For example, for “Sylvia Costa” in the script “Death Wears White”, it is:
@@ D. Father Tom is hate of Drake Li Morette
Table 4: Examples of Each Type of Our Evaluation Questions
@@ 10 Do you have any personal opinions or theories about the case?
Table 5: Predefined questions for Werewolf method.
@@ B.2 Evaluation questions
The evaluation questions are devised based on our annotations and aim to assess the understanding of the intricate relationships and narratives present within the game scripts. These questions are segmented into three categories, with each category of questions and examples shown in Table 4.
@@
B.3 Annotation Details
Due to the cost of human annotation on such a long and complicated corpus, we believe our current dataset is sufficient for evaluation purposes. Our dataset can serve as a starting point to evaluate various strategies and boost investment in larger-scale labelling and training for agents with enhanced inferential capabilities.
B.4 Disclaimers of Risks
B.5 Human Subject Evaluation Setup
Since the same player would know the killer’s identity if they tried different models, we recruited four players to act as the murderer, each playing through three scripts (Sin, Deadly Fountain, Riverside Inn) across four methods, totalling 49 hours of gameplay. Each human player assumed the role of one character in the game, while the other characters were controlled by Questum -powered agents.
Evaluation Werewolf O-CoT ThinkThrice Questum
Story Advancement $1.50_{\pm 0.37}$ $2.08_{\pm 0.14}$ $2.67_{\pm 0.00}$ $3.50_{\pm 0.17}$
Question Quality $1.67_{\pm 0.41}$ $1.58_{\pm 0.36}$ $2.08_{\pm 0.28}$ $3.25_{\pm 0.28}$
Response Quality $2.83_{\pm 0.17}$ $2.92_{\pm 0.36}$ $3.00_{\pm 0.33}$ $3.50_{\pm 0.29}$
Response Speed $1.17_{\pm 0.17}$ $1.33_{\pm 0.41}$ $1.17_{\pm 0.17}$ $2.50_{\pm 0.17}$
Role Immersion $2.25_{\pm 0.36}$ $2.17_{\pm 0.5}$ $2.33_{\pm 0.24}$ $2.83_{\pm 0.29}$
overall $1.88_{\pm 0.59}$ $2.01_{\pm 0.55}$ $2.25_{\pm 0.62}$ $3.11_{\pm 0.39}$
Table 6: Results of Agent-vs-Human Evaluation using Human-Centric metrics.
To ensure fairness, the human players were given the same information and rules as the agents, maintaining consistent conditions across both evaluation scenarios. Both the human players and the agents shared the same objectives: to achieve victory in the game by fulfilling their individual objectives. This setup allowed for a direct comparison between human-agent and agent-agent interactions, providing insights into Questum ’s ability to adapt and perform in mixed environments.
Score Story Advancement Question Quality Response Quality Role Immersion Response Speed
5 points Actively gathers information and conducts in-depth analyses. Reasoning is rigorous, offering key insights that significantly advance the narrative. Poses highly relevant, in-depth questions that directly propel the plot and uncover hidden clues. Provides accurate and comprehensive responses fully aligned with the character’s background and narrative needs. Offers clear, valuable information. Fully embodies the character. Actions, dialogue, and emotional expressions are authentically aligned with the character’s traits, leaving a strong impression. Responds promptly with no perceptible delay. The thought process is fluid and maintains a brisk, engaging pace.
4 points Proactively seeks information and performs reasonable analysis. Reasoning is generally accurate and supports narrative progression. Most questions are relevant and yield useful information. Occasionally introduces innovative queries. Generally provides accurate and sufficiently complete answers. Minor omissions occur but do not hinder overall understanding. Frequently demonstrates character traits. Occasionally shows minor deviations, but emotional expression is largely appropriate. Responds relatively quickly with brief, occasional delays. The thought process is mostly smooth and well-paced.
3 points Participates in information gathering but lacks analytical depth. Reasoning may be somewhat biased, offering moderate narrative progression. Some questions are relevant, but lack depth. Occasional repetition or irrelevant queries occur. Responses are basically accurate but sometimes incomplete or vague. May require further clarification by others. Occasionally exhibits character traits, but inconsistently. Emotional expression is average, and immersion is limited. Moderate response speed with some noticeable pauses. The reasoning process is acceptable but not seamless.
2 points Shows limited initiative in gathering information, with superficial analysis. Reasoning is frequently flawed, hindering effective plot advancement. Most questions bear little relevance, often fail to obtain valuable clues, and frequently repeat similar inquiries. Responses are often inaccurate, incomplete, or misleading. Frequent clarification requests from others are necessary. Characterization is unclear; actions and dialogue often stray significantly from the intended character profile. Emotional expression is awkward or absent. Responses are slow with frequent delays. The reasoning process is disjointed, disturbing the overall gameplay flow.
1 point Does not gather information or attempt analysis. Unable to perform meaningful reasoning, leaving the narrative stagnant. Poses irrelevant or no questions, providing no support for narrative progression. Responses are unrelated or outright refusals. Offers no assistance in understanding the narrative. Demonstrates no character immersion; actions and dialogue contradict character traits. No emotional expression is evident. Extremely slow responses or prolonged silence. The reasoning process halts, severely impeding the game’s progress.
Table 7: Criteria for each evaluation dimension (Story Advancement, Question Quality, Response Quality, Role Immersion, and Response Speed).
Appendix C Implementation
C.1 Implementation Details
For retrieval enhancement, we utilized the FAISS [^4] https://github.com/facebookresearch/faiss library to construct a vector database, creating FAISS indices using the L2 distance metric. Embeddings were generated using the Azure API’s text-embedding-ada-002 service. Scripts were divided into segments, with each segment having a maximum length of 50 tokens. For dialogue records, a question-and-answer pair was stored as a single segment. During retrieval, the maximum script and dialogue lengths included in the prompt were set to 4000 tokens. For evaluation, these maximum lengths were increased to 5000 tokens. Additionally, the hyperparameters were set with $\epsilon$ equal to 0.1 and $\beta$ equal to 0.2.
Experiment
C.2 Overall Performance Computing
@@
C.3 Comparison Models
@@
      Selecting questions from a predefined question written by the human specialist, based on the current dialogue and script. The list of expertly devised questions is presented in Table 5.